# GIST: Greedy Independent Set Thresholding for Diverse Data Summarization

### v2

@@ 4.1 General Metric Spaces
We set $d$ to be $(1-\nicefrac{{1}}{{e}})U$ where $U$ is the cardinality of the union of all sets, i.e., $U=|\bigcup_{S\in\mathcal{F}}S|$. In the YES case, the optimum solution is the family of $k$ disjoint sets from different groups that cover the entire universe. The value of optimum in this case is $2d+U=(2(1-\nicefrac{{1}}{{e}})+1)U$.
@@
Since $k-1$ is at least $\nicefrac{{1}}{{\varepsilon}}$, the submodular value in case (a) does not exceed:
@@
4.2 Linear Utility
For the special case of linear utility functions, we prove a tight hardness-of-approximation result complementing the $\nicefrac{{2}}{{3}}$-approximation guarantee in Theorem 3.4.
Theorem 4.2.
[…]
We give the proof in Section C.1. Our construction builds on the work of Håstad (1996); Zuckerman (2006), which shows that the maximum clique problem does not admit an efficient $n^{1-\theta}$-approximation algorithm, for any constant $\theta>0$.
4.3 Euclidean Metric
[…]
Lemma 4.3 (Alimonti and Kann (2000, Theorem 3.2)).
Our second ingredient is a node embedding function $h_{G}(v)$ that encodes graph adjacency in Euclidean space.
Lemma 4.4.
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\frac{1}{2(\Delta+1)},$
and if $\{u,v\}\not\in E$ then $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
Theorem 4.5.
The MDMS problem is APX-complete for the Euclidean metric even if $g$ is a linear function.
We sketch the main idea below and defer complete proofs to Appendix C. Let $\mathcal{I}(G)$ be the set of independent sets of $G$. Using Lemma 4.4, for any set of nodes $S\subseteq V$ with $|S|\geq 2$ in a graph $G$ with max degree $\Delta\leq 3$, we have the property:
  • $S\in\mathcal{I}(G)\implies\texttt{div}(S)=1$;
The gap between the cases is at least $\nicefrac{{1}}{{8}}$ for all such graphs (i.e., a universal constant), allowing us to prove that MDMS inherits the APX-completeness of bounded-degree max independent set.
5 Experiments
5.1 Synthetic Dataset
We use a collection of weighted random points and a budget-additive (submodular) utility function to (1) compare against baselines and (2) highlight the non-monotone behavior.
Setup.
Generate $n=1000$ points $\mathbf{x}_{i}\in\mathbb{R}^{d}$, for $d=64$, where each $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ is i.i.d. and assigned a uniform continuous weight $w_{i}\sim\mathcal{U}_{[0,1]}$. For $\alpha,\beta\in[0,1]$, the objective function below trades off between the average (capped) utility of the selected points and their min-distance diversity:
We consider three baselines: random, simple, and greedy. random chooses a random set of $k$ points, permutes them, and returns the best prefix since the objective function is non-monotone. simple is the $0.387$-approximation algorithm in Section 2.1. greedy builds $S$ one point at a time by selecting the point with index $i^{*}_{t}=\operatorname*{arg\,max}_{i\in V\setminus S_{t-1}}f(S_{t-1}\cup\{i\})-f(S_{t-1})$ at each step $t\in[k]$, and outputs the best prefix $S=\operatorname*{arg\,max}_{t\in[k]}f(S_{t})$.
Results.
We plot the values of $f(S)$ for the baseline methods and GIST for all $k\in[n]$ in Figure 1. We set $\alpha=0.95$ to roughly balance the submodular and diversity terms. [^3] If $\mathbf{x},\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ are i.i.d., then $\mathbb{E}[\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}]=2d$. Setting $\frac{\alpha}{2}\approx(1-\alpha)\sqrt{2d}\implies\alpha\approx 1-\frac{1}{2\sqrt{2d}+1}=0.957$, for $\beta\geq 0.5$. We sweep over $\alpha$ and budget $\beta$ in Section D.1, and plot how these hyperparameters change $f(S)$. Note that the objective values are not directly comparable as a function of $k$ because of the $\nicefrac{{1}}{{k}}$ normalizing factor (hence the discontinuities), but both terms in the objective are decreasing on average in $k$ for a fixed $\alpha$. GIST dominates simple, greedy, and random in all cases. The poor performance of greedy is somewhat surprising since it is often competitive for subset selection tasks. simple beats greedy for all but the extreme values of $k$, i.e., small or close to $n$, but is always worse than GIST.
Figure 1: Baseline comparison plotting $f(S_{\text{ALG}})$ for $k\in[n]$ on synthetic data with $n=1000$, $\alpha=0.95$, and $\beta=0.75$.
5.2 Image Classification
Setup.
We use the standard vision dataset ImageNet (Russakovsky et al., 2015) containing ~1.3 million images and 1000 classes. We select 10% of the points at random and use them to train an initial ResNet-56 model $\bm{\theta}_{0}$ (He et al., 2016). Then we use this initial model to compute an uncertainty value and 2048-dimensional embedding for each example. The margin-based uncertainty value of $\mathbf{x}_{i}$ is given by $u_{i}=1-(\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})-\Pr(y=b^{\prime}\mid\mathbf{x}_{i};~{}\bm{\theta}_{0}))$, which measure the difference between the probability of the best and second-best class label $b^{\prime}$ for an example. Finally, we use the fast maximum inner product search of Guo et al. (2020) to build a $\Delta$-nearest neighbor graph in the embedding space using $\Delta=10$ and cosine distance. We present all model training hyperparameters in Appendix D.
We compare GIST with several state-of-the-art benchmarks:
  • random: We select samples from the dataset uniformly at random (without replacement). While this approach is simple and lightweight, it creates diversity in many settings and provides good solutions.
  • margin (Roth and Small, 2006): Margin sampling selects the top-$k$ points based on how hard they are to classify, i.e., using margin-based uncertainty values $u_{i}$.
  • $k$-center (Sener and Savarese, 2018): We run the classic greedy algorithm for $k$-center on the $\Delta$-nearest neighbor graph. The distance between non-adjacent nodes is implicitly assumed to be very large, which is reasonable in the large budget regime $k=\Omega(n)$.
  • submod: We select a subset by maximizing the submodular objective function $\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j)$, under the budget constraint $|S|\leq k$. Similar pairwise submodular objective functions have been used in (Iyer et al., 2021; Kothawade et al., 2021; Ramalingam et al., 2021). Here $s(i,j)$ denotes the similarity between adjacent nodes using the same graph as in $k$-center. We tuned for best performance by selecting $\alpha_{s}=0.9$ and $\beta_{s}=0.1$.
We evaluate GIST using both linear and submodular utility functions. Let $f(S)=\alpha\cdot g(S)+(1-\alpha)\cdot\texttt{div}(S)$ for $\alpha\in[0,1]$. In Table 1, GIST-linear uses the same linear utility as margin: $g(S)=\sum_{i\in S}u_{i}$ with $\alpha=0.9$. GIST uses the same submodular utility as submod: $g(S)=\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j)$ with $\alpha=0.95$, $\alpha_{s}=0.9$, and $\beta_{s}=0.1$. We use the same graph as $k$-center for distances.
Table 1: Top-1 classification accuracy (%) on ImageNet for different single-shot data downsampling algorithms. The cardinality constraint $k$ is expressed as a fraction of the ~1.3 million examples. The results are the average of 3 trials and the top performance for each $k$ is shown in bold.
$k$ (of data) random margin $k$-center submod GIST-linear GIST
30 66.27 65.97 66.70 66.52 66.89 67.23
40 69.17 69.73 70.06 70.71 70.51 70.75
50 71.05 72.33 73.09 72.69 72.67 73.14
60 72.49 73.42 73.60 74.20 74.37 74.30
70 73.70 74.49 74.24 75.24 75.41 75.32
80 74.42 75.01 75.17 75.91 75.96 75.45
90 75.16 75.10 74.97 75.84 76.12 76.03
Results.
We run each sampling algorithm with cardinality constraint $k$ on the full dataset to get a subset of examples that we then use to train a new ResNet-56 model. We report the top-1 classification accuracy of these models in Table 1. GIST with a linear or submodular utility is superior to all baselines. Interestingly, there is a cutover value of $k$ where GIST-linear becomes the best algorithm. Finally, observe that GIST-linear outperforms margin in all cases. This demonstrates how $\texttt{div}(S)$ encourages diversity in the subset of sampled points and ultimately improves downstream model quality.
Conclusion
This work presents a novel subset selection problem called MDMS that combines the total utility of the selected points (modeled by a monotone submodular function $g$) with the $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ diversity objective. We design and analyze the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee by solving a series of maximum independent set instances on geometric intersection graphs with the GreedyIndependentSet bicriteria approximation algorithm. We complement our algorithm with a $0.5584$-hardness of approximation. From the theoretical perspective, it is an interesting open problem to bridge the gap between the algorithmic and hardness results of $0.5$ and $0.5584$ approximations.
For the special case of a linear utility function $g$, we show that GIST has a $\nicefrac{{2}}{{3}}$-approximation guarantee, and we complement this with a $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$.
Finally, we present an empirical study showing that GIST steadily outperforms baseline methods for MDMS (in particular greedy) on a simple set of Gaussian points with random weights. We conclude by comparing the top-1 image classification accuracy achieved by GIST and three other state-of-the-art data sampling methods for ImageNet, demonstrating the benefit of optimizing for a blend of utility and diversity.
Appendix A $f(S)$ is Not Submodular
Consider the instance on four points $N=\{a,b,c,d\}\subseteq\mathbb{R}$ where $a=0$, $b=1$, and $c=d=2$ in the (one-dimensional) Euclidean metric, i.e., all the points are collinear. Let $g(S)=0$ for all $S\subseteq V$. For $f(S)$ to be submodular, it must hold that for every $S\subseteq T\subseteq N$ and $x\in N\setminus T$,
$f(S\cup\{x\})-f(S)\geq f(T\cup\{x\})-f(T).$
However, if $x=b$, $S=\{a,c\}$ and $T=\{a,c,d\}$, we have:
$f(S\cup\{x\})-f(S)$ $=f(\{a,b,c\})-f(\{a,c\})=1-2=-1$
$f(T\cup\{x\})-f(T)$ $=f(\{a,b,c,d\})-f(\{a,c,d\})=0-0=0,$
so $f(S)$ is not submodular.
In this instance, the function $f$ is not monotone. However, a similar monotone but still not submodular $f$ can be defined by setting $g(S)=\sum_{v\in S}w(v)$ where $w(v)=2$ for every $v\in N$.
Appendix B Proof of Theorem 3.4
Lemma B.1.
[…]
Proof.
$<d^{\prime}+d^{\prime}$
$\leq d/2+d/2$
$=d,$
Now let $C_{i}\subseteq V$ be the set of uncovered (by the open balls) points that become covered when GreedyIndependentSet selects $t_{i}$. Concretely, $C_{i}=B_{i}\setminus(B_{1}\cup\cdots\cup B_{i-1})$. Each $C_{i}$ contains at most one point in $S_{d}^{*}$ since $|B_{i}\cap S_{d}^{*}|\leq 1$. Moreover, if $s\in C_{i}\cap S_{d}^{*}$, then $w(t_{i})\geq w(s)$ because the points are sorted in non-increasing order and selected if uncovered.
It remains to account for the points in $S_{d}^{*}\setminus A$. Note that if we have any such points, then $|T|=k$ since the points in $S_{d}^{*}\setminus A$ are uncovered at the end of the algorithm. Further, for any $t_{i}\in T$ and $s\in S_{d}^{*}\setminus A$, we have $w(t_{i})\geq w(s)$ since $t_{i}$ was selected and the points are sorted by non-increasing weight. Therefore, we can assign each $s\in S_{d}^{*}\setminus A$ to a unique $C_{i}$ such that $C_{i}\cap S_{d}^{*}=\varnothing$. It follows that
Adding the two sums together in Equation 3 and (4) completes the proof. ∎
See 3.4
Proof of Theorem 3.4.
Let $d^{*}$ be the minimum distance between two distinct points in $S^{*}$. There are two cases: $d^{*}<\varepsilon d_{\max}$ and $d^{*}\geq\varepsilon d_{\max}$. In the first case, outputting the $k$ heaviest points (Line 2) achieves a $(1-\varepsilon)$-approximation. To see this, first observe that
$\textnormal{OPT}\geq d_{\max}>\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}>d^{*}.$
The sum of the $k$ heaviest points upper bounds $g(S^{*})$, so we have
$\textnormal{ALG}\geq g(S^{*})=\textnormal{OPT}-d^{*}>(1-\varepsilon)\textnormal{OPT}.$
Therefore, we focus on the second case where $d^{*}\geq\varepsilon d_{\max}$. GIST considers a threshold $d=(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2\in D$ such that
$\frac{d^{*}}{2}\in[d,(1+\varepsilon)d)\implies\frac{1}{1+\varepsilon}\cdot\frac{d^{*}}{2}<d\leq\frac{d^{*}}{2}.$
[…]
$f(T)$ $\geq g(S^{*})+d\geq g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}.$ (5)
The potential max-diameter update to $S$ on Line 5 gives us
$\textnormal{ALG}\geq d_{\max}\geq d^{*}.$ (6)
Putting (5) and (6) together, for any $0\leq p\leq 1$, it holds that
ALG $\geq p\cdot\left(g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}\right)+(1-p)\cdot d^{*}$
$=p\cdot g(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot d^{*}.$
To maximize the approximation ratio as $\varepsilon\rightarrow 0$, we set $p=2/3$ by solving $p=1-p/2$. It follows that
ALG $\geq\frac{2}{3}\cdot g(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot d^{*}$ (7)
$=\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot d^{*}$
$\geq\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot d^{*}$
$\geq\left(\nicefrac{{2}}{{3}}-\varepsilon\right)\cdot\textnormal{OPT},$
which completes the proof. ∎
Appendix C Missing Analysis for Section 4
C.1 Proof of Theorem: 4.2
@@ Appendix C Missing Analysis for Section 4
[…]
@@
Use the same threshold value of $k$ (in the YES and NO instance above) for the cardinality constraint on set $S$, and set each weight $w(v)=1$. In a YES instance, selecting a clique of size $k$ as set $S$ results in the maximum possible value of the objective:
@@
These two terms become equal for $\alpha=1/(2-\delta)$. Thus, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is
@@
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$ $=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
@@
[…]
@@
We reduce this instance of bounded-degree maximum independent set to MDMS with objective function (2) as follows. Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma 4.4. We use the same threshold value of $k$ (between YES and NO instances above) for the cardinality constraint on set $S$, and we set each weight $w(v)=1$. In a YES instance, selecting an independent set of size $k$ as the set $S$ results in the maximum value of objective (2):
@@
In a NO instance, the best objective value that can be achieved in polynomial-time is the maximum of the following two scenarios: (a) selecting $k$ points with minimum distance at most $81-1/(2(\Delta+1))=1-1/$ 8, or (b) selecting at most $(1-\varepsilon_{0})k$ vertices forming an independent set with minimum distance equal to $1$. The maximum value obtained by any polynomial-time algorithm is then
@@
These two terms become equal for $\alpha=1/(1+8\varepsilon_{0})$. Therefore, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is upper bounded by
@@
Appendix D Additional Details for Section 5
D.1 Synthetic Dataset
We extend our comparison of baseline algorithms in Section 5.1 by sweeping over values of $\alpha,\beta$ in the objective function.
@@ D.2 Image Classification
We generate predictions and embeddings for all points using a coarsely-trained ResNet-56 model (He et al., 2016) trained on a random 10% subset of ImageNet (Russakovsky et al., 2015). We use SGD with Nesterov momentum 0.9 with 450/90 epochs. The base learning rate is 0.1, and is reduced by a tenth at 5, 30, 69, and 80. We extract the penultimate layer features to produce $2048$-dimensional embeddings of each image. We use the same hyperparameters as the original ResNet paper (He et al., 2016) with budgets and one-shot subset selection experiments designed in the same manner as Ramalingam et al. (2021).

### v3

@@ Appendix D Missing analysis for Section 4
We set $d$ to be $(1-\nicefrac{{1}}{{e}})U$ where $U$ is the cardinality of the union of all sets, i.e., $U=|\bigcup_{S\in\mathcal{F}}S|$. Finally, we set $\lambda=1$ to complete the construction of the MDMS instance.
In the YES case, the optimum solution is the family of $k$ disjoint sets from different groups that cover the entire universe. The value of optimum in this case is $2d+U=(2(1-\nicefrac{{1}}{{e}})+1)U$.
@@
Since k1k-1 is at least $\nicefrac{{1}}{{\varepsilon}}$, the submodular value in case (a) does not exceed:
@@
D.2 Proof of Theorem 4.2
@@
First, recall that a clique is a subset of vertices in an undirected graph such that there is an edge between every pair of its vertices. Håstad [28] and Zuckerman [64] showed that the maximum clique problem does not admit an $n^{1-\theta}$-approximation for any constant $\theta>0$, unless $\textnormal{NP}=\textnormal{P}$. This implies that there is no constant-factor approximation algorithm for maximum clique. In other words, for any constant $0<\delta\leq 1$, there exists a graph $G$ and a threshold integer value $k$ such that it is NP-hard to distinguish between the following two cases:
@@
Use the same threshold value of $k$ (in the YES and NO instance above) for the cardinality constraint on set $S$, and set each weight $w(v)=\nicefrac{{\alpha}}{{k}}$ for some parameter $\alpha$ that we set later in the proof. We also set $\lambda=1-\alpha$. In a YES instance, selecting a clique of size $k$ as set $S$ results in the maximum possible value of the objective:
@@
We make these two terms equal by setting $\alpha=1/(2-\delta)$. Thus, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is
@@
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$
$=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
@@
We build on the hardness of approximation for the maximum independent set problem for graphs with maximum degree $\Delta=3$. Alimonti and Kann [4, Theorem 3.2] showed that this problem is APX-complete, so there exists an $\varepsilon_{0}>0$ such that there is no polynomial-time $(1-\varepsilon_{0})$-approximation algorithm unless $\textnormal{NP}=\textnormal{P}$. Hence, there exists a graph $G$ with max degree $\Delta=3$ and a threshold integer value $k$ such that it is NP-hard to distinguish between the following two cases:
@@
We reduce this instance of bounded-degree maximum independent set to MDMS with objective function (2) as follows. Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma˜4.4. We use the same threshold value of $k$ (between YES and NO instances above) for the cardinality constraint on set $S$, and we set each weight $w(v)=\nicefrac{{\alpha}}{{k}}$ for some parameter $\alpha$ that we set later in the proof. We also set $\lambda=1-\alpha$.
In a YES instance, selecting an independent set of size $k$ as the set $S$ results in the maximum value of objective (2):
@@
In a NO instance, the best objective value that can be achieved in polynomial-time is the maximum of the following two scenarios: (a) selecting $k$ points with minimum distance at most $1-1/(2(\Delta+1))=1-1/8$, or (b) selecting at most $(1-\varepsilon_{0})k$ vertices forming an independent set with minimum distance equal to $1$. The maximum value obtained by any polynomial-time algorithm is then
@@
We make these two terms equal by setting $\alpha=1/(1+8\varepsilon_{0})$. Therefore, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is upper bounded by
@@
Appendix E Additional details for Section 5
E.1 Synthetic dataset
We extend our comparison of baseline algorithms in Section˜5.1 by sweeping over values of $\alpha,\beta$ in the objective function.
@@ E.2 Image classification
We generate predictions and embeddings for all points using a coarsely-trained ResNet-56 model [29] trained on a random 10% subset of ImageNet [54]. We use SGD with Nesterov momentum 0.9 with 450/90 epochs. The base learning rate is 0.1, and is reduced by a tenth at 5, 30, 69, and 80. We extract the penultimate layer features to produce $2048$-dimensional embeddings of each image. We use the same hyperparameters as the original ResNet paper [29] with budgets and one-shot subset selection experiments designed in the same manner as [49].
Running times.
The end-to-end running time is dominated by ImageNet model training, which takes more than a few hours even with accelerators (e.g., GPU/TPU chips). The subset selection algorithms that use margin and submodular sampling range between 3–4 minutes per run on an average. GIST subset selection is similar to the margin or submodular algorithms for a given distance threshold $d$. By using parallelism for different $d$ values, we can keep the GIST-submod subset selection runtime the same as the submod algorithm. In summary, the actual subset selection algorithm step is extremely fast (nearly negligible) compared to the ImageNet training time. Furthermore, we can exploit distributed submodular subset selection algorithms that can even handle billions of data points efficiently [15].